OpenStudy (anonymous):

Help!!: Write the equation of a line that is parallel to the line connecting (2,5) and (-1,-4)

OpenStudy (sleepyhead314):

Hi there @angelmaxine and Welcome to OpenStudy! :) Sorry for getting here a little late, I can try to help you with your problem:

OpenStudy (sleepyhead314):

An equation of a line that is parallel to another line would have the Same Slopes sooo we should first try to find the slope between the points (2, 5) and (-1, -4)

OpenStudy (anonymous):

@sleepyhead314 it's okay, and thank you. I'm not sure how to maneuver open study. But I need help with pre calculus!/:

OpenStudy (anonymous):

So we would use y1-y2/x1-x2?

OpenStudy (sleepyhead314):

yes! :) that's right what do you think we would get?

OpenStudy (sleepyhead314):

it is ok that you might not know how to get around the site a friend of mine created a tutorial on how to use the site that you can check out sometime ^_^ http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

OpenStudy (sleepyhead314):

\[slope = \frac{ y_2 - y_1 }{ x_2 - x_1 } = \frac{-4 - 5}{-1 - 2}\]

OpenStudy (anonymous):

So.. Positive 3??

OpenStudy (sleepyhead314):

that's perfect! :) great job!! then we can plug that into the equation y = mx + b

OpenStudy (anonymous):

That's where I keep getting stuck.

OpenStudy (sleepyhead314):

that's ok :) we know that m = slope and b = y-intercept right?

OpenStudy (sleepyhead314):

in this question "b" can be anything because it just wants the line to be parallel

OpenStudy (sleepyhead314):

does this make sense @angelmaxine ? :)

OpenStudy (anonymous):

No :/

OpenStudy (sleepyhead314):

ok, what part do you not get? I'll try to explain better :)

OpenStudy (anonymous):

Do I just pick one set ? Like (2,5) or (-1,-4) to plug in?

OpenStudy (sleepyhead314):

we would plug in the slope = 3 that we got for "m" so y = 3x + b ^ right?

OpenStudy (sleepyhead314):

then "b" can be any number you want it to be for this question :)

OpenStudy (sleepyhead314):

for example y = 3x + 1 would work y = 3x + 7 would also work :P

OpenStudy (sleepyhead314):

this works because all they wanted was the equation of a line that is Parallel and all a line needs to be Parallel is the same slope that's why any "b" would work :)

OpenStudy (sleepyhead314):

so your answer can be as simple as y = 3x + 1

OpenStudy (anonymous):

How'd you get 7??